OpenStudy (anonymous):

Ms. Rao buys a computer, a printer, and a scanner for $2,543. The computer cost $1,502 more then the printer. The printer cost $123 more than the scanner. How much does the computer cost? can someone please help me?

OpenStudy (anonymous):

Ok we know the scanner is 2543 and that the printer is 123 more than that: So the printer price is 2543 + 123 = $2666 We are given that the computer costs 1502 more than the printer So we do 1502 + 2666 = $4168

OpenStudy (anonymous):

The following equations were derived from the problem text.\[\{c+p+s=2543,c-1502=p,p-123=s\} \]

OpenStudy (anonymous):

scanner =x printer = x+123 computer = x+123+1503 2453 =scanner + pinter +computer . x+x+123+x+123+1503 = 2543 find x that will be price of scanner and subsitute this x in above equations to find price of computer and printer

OpenStudy (anonymous):

@AustinC Mathematica says that that the solution to x+x+123+x+123+1503 = 2543 is\[\frac{794}{3} \]One would expect an integer solution.

OpenStudy (anonymous):

Mathematica can calculate the problem solution in one line using it's function, Eliminate or it's Solve function. Refer to the attachment.

OpenStudy (anonymous):

Wow, thats pretty cool